# AI-Powered Inverse Design of Ku-Band SIW Resonant Structures by Iterative Residual Correction Network

### v1

@@ 4.3 Practical Validation through Simulation and Fabrication
(a) Structure with 3 in-band resonances.
@@
(b) Structure with 4 in-band resonances.
@@
Figure 13: Comparison between the measured and simulated S-parameters for the proposed multi-mode SIW structures: (a) structure with 3 in-band resonances, (b) structure with 4 in-band resonances.
Figure 14: Measurement setup for the fabricated SIW resonator. The structure is mounted between two waveguide-to-coaxial adapters for S-parameter characterization.
@@ 5 Conclusion
In this work, we presented an efficient and accurate inverse design framework for Ku-band SIW components based on multimode resonators using a novel deep learning architecture named the IRC-Net. The proposed method integrates an initial FIM with a residual-based correction mechanism inspired by the HiFR2-Net framework, resulting in significantly enhanced prediction accuracy and generalization capabilities. Using multimode resonance structures, we demonstrated the feasibility of controlling resonance frequencies, which enables the design of complex and compact SIW-based filters and resonant components. To validate our approach, we first designed and fabricated a three-resonance SIW structure using conventional full-wave simulations, which was then used to train the IRC-Net model. The trained model was subsequently used to predict the geometric parameters of a more complex four-resonance SIW structure based solely on the desired frequency response. Both predicted structures were fabricated and tested experimentally. The measured S parameters closely matched the simulated and predicted responses, confirming the practical viability of the proposed model. Compared to traditional inverse design methods, IRC-Net significantly reduces prediction error and improves consistency across a wide range of frequency responses, as demonstrated by comprehensive statistical evaluations. This study not only introduces a new learning-based methodology for inverse electromagnetic modeling but also establishes a foundation for smart filter synthesis using data-driven techniques. The IRC-Net framework offers a scalable and generalizable solution for the inverse design of high-performance microwave components, paving the way for future research in intelligent and adaptive RF/microwave systems.

### v2

@@ 4.3 Parametric Evaluation and Experimental Validation
(a) Measured vs. predicted S-parameters for the structure with three in-band resonances.
@@
(b) Measured vs. simulated S-parameters for the structure with four in-band resonances.
@@
Figure 13: Comparison between the measured, simulated and predicted S-parameters for the fabricated SIW structures: (a) three-resonance structure, (b) four-resonance structure.
@@ 5 Conclusion
We have presented a three-stage progressive deep learning framework for the inverse design of multi‑mode SIW filters, comprising a Feedforward Inverse Model, a  Hybrid  Inverse‑Forward  Residual  Refinement  Network, and an Iterative Residual Correction Network (IRC‑Net). Experimental validation on fabricated prototypes—including geometries outside the training set—demonstrated that IRC‑Net consistently outperforms the baseline models, reducing mean squared error from 0.00191 to 0.00146 and mean absolute error from 0.0262 to 0.0209, while enabling sub‑millisecond prediction times. This approach significantly reduces dependence on computationally intensive electromagnetic simulations and facilitates rapid prototyping of complex microwave components.
Although the demonstrated results focus on symmetric filters with up to four resonators, the modular and iterative architecture of IRC‑Net holds promise for extension to higher-order and asymmetric SIW configurations, as well as broader applications across microwave and millimeter-wave domains. However, two main limitations remain. First, the generalization capability of the model across more diverse structures—such as those with varying topology, material properties, or irregular geometries—has not yet been fully explored. Future work should investigate training on larger and more heterogeneous datasets to improve robustness and applicability to a wider range of practical designs.
Second, like most data-driven models, the proposed framework operates as a black box, offering limited insight into the underlying physical reasoning behind its predictions. To address this, future research may incorporate concepts from explainable artificial intelligence (XAI) to improve interpretability and trustworthiness, potentially revealing new physics-informed design principles embedded within the network.
CRediT authorship contribution statement
Mohammad Mashayekhi: Conceptualization, Methodology, Software, Data curation, Project administration, Writing – original draft. Kamran Salehian: Conceptualization, SIW structure design, Fabrication of SIW, Figure preparation for SIW design, Writing – review & editing. Abbas Ozgoli: Fabrication and measurement of SIW. Saeed Abdollahi: Writing – introduction. Abdolali Abdipour: Supervision. Ahmed A. Kishk: Supervision, Project administration, Writing – review & editing.
Funding
This work is self-funded research. NO Government, External organizations are involved with financial assistance.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
The dataset generated and analyzed during the current study is not publicly available due to its integration into ongoing research projects and future planned publications.